OpenStudy (anonymous):

Verify the trig equation by substituting identities to match the right side of the equation to the left side. -tan^2x + sec^2x = 1

OpenStudy (campbell_st):

well tan= sin/cos and sec = 1/cos so multiply every term by cos^2 \[-\frac{\sin^2(x)}{\cos^2(x)} \times \cos^2(x) + \frac{1}{\cos^2(x)} \times \cos^2(x) = 1 \times \cos^2(x)\] do the multiplication and rewrite the equation for the solution.

OpenStudy (anonymous):

First note the following:\[\bf \tan^2(x)+1=\sec^2(x)\]Substituting for sec^2(x) we get:\[\bf -\tan^2(x)+(\tan^2(x)+1)=1\]Opening up the brackets we get:\[\bf -\tan^2(x)+\tan^2(x)+1=1\]From here it's clear to see that the L.H.S = R.H.S. Can you complete the proof however by adding a few more steps? @FLBaby123

OpenStudy (anonymous):

Can you explain how to multiply them? Because i'm super confused. @campbell_st

OpenStudy (campbell_st):

ok... well by mutliplying you will cancel the denominator